\newpage

\begin{thebibliography}{99}

\bibitem{horava}
P. Ho\v{r}ava, ``Quantum gravity at a Lifshitz point'', Phys. Rev. D79 (2009) 084008, arXiv:0901.3775.

\bibitem{scalarvector1}
P. Ho\v{r}ava, ``Quantum criticality and Yang-Mills gauge theory'', arXiv:0811.2217;
B. Chen, Q. G. Huang, ``Field theory at a Lifshitz point'', Phys. Lett. B683 (2010) 108, arXiv:0904.4565.

\bibitem{scalar1}
G. Calcagni, ``Cosmology of the Lifshitz universe'', JHEP 0909 (2009) 112, arXiv:0904.0829.

\bibitem{scalar2}
G. Calcagni, ``Detailed balance in Ho\v{r}ava-Lifshitz gravity'', Phys. Rev. D81 (2010) 044006, arXiv:0905.3740.

\bibitem{scalarvector2}
E. Kiritsis, G. Kofinas, ``Ho\v{r}ava-Lifshitz cosmology'', Nucl. Phys. B821 (2009) 467, arXiv:0904.1334.

\bibitem{magnetic}
S. Maeda, S. Mukohyama and T. Shiromizu, ``Primordial magnetic field from non-inflationary cosmic expansion in Ho\v{r}ava-Lifshitz gravity'', Phys. Rev. D80 (2009) 123538, arXiv:0909.2149.

\bibitem{vector}
J. M. Romero, J. A. Santiago, O. Gonz\'{a}lez-Gaxiola and A. Zamora ``Electrodynamics \`{a} la Ho\v{r}ava'', Mod. Phys. Lett. A25 (2010) 3381, arXiv:1006.0956.

\bibitem{problems}
I. Kimpton and A. Padilla, ``Lessons from the decoupling limit of Ho\v{r}ava gravity'', JHEP 1007 (2010) 014, arXiv:1003.5666;
J. Bellorin and A. Restuccia, ``On the consistency of the Ho\v{r}ava theory'', arXiv:1004.0055;
D. Blas, O. Pujolas and S. Sibiryakov, ``Models of non-relativistic quantum gravity: the good, the bad and the healthy'', JHEP 1104 (2011) 018, arXiv:1007.3503;
A. Padilla, ``The good, the bad and the ugly .... of Ho\v{r}ava gravity'', J. Phys. Conf. Ser. 259 (2010) 012033, arXiv:1009.4074;
T. P. Sotiriou, ``Ho\v{r}ava-Lifshitz gravity: a status report'', J. Phys. Conf. Ser. 283 (2011) 012034, arXiv:1010.3218.

\bibitem{CMP}
C. G. Callan, R. C. Myers and M. J. Perry, ``Black holes in string theory'', Nucl. Phys. B311 (1989) 673, http://ccdb4fs.kek.jp/cgi-bin/img/allpdf?200036442

\bibitem{deformed}
H. Lu, J. Mei and C. N. Pope, ``Solutions to Ho\v{r}ava gravity'', Phys. Rev. Lett. 103 (2009) 091301, arXiv:0904.1595;
H. Nastase, ``On IR solutions in Ho\v{r}ava gravity theories'', arXiv:0904.3604;
A. Kehagias and K. Sfetsos, ``The black hole and FRW geometries of non-relativistic gravity'' Phys. Lett. B678 (2009) 123, arXiv:0905.0477;
M. I. Park,``The black hole and cosmological solutions in IR modified Ho\v{r}ava gravity'', JHEP 0909 (2009) 123, arXiv:0905.4480.

\bibitem{charged}
R. G. Cai, L. M. Cao and N. Ohta, ``Topological black holes in Ho\v{r}ava-Lifshitz gravity'', Phys. Rev. D80 (2009) 024003, arXiv:0904.3670;
E. \'{O}. Colg\'{a}in and H. Yavartanoo, ``Dyonic solution of Ho\v{r}ava-Lifshitz gravity'', JHEP 0908 (2009) 021, arXiv:0904.4357;
S. S. Kim, T. Kim and Y. Kim, ``Surplus solid angle as an imprint of Ho\v{r}ava-Lifshitz gravity'', Phys. Rev. D80 (2009) 124002, arXiv:0907.3093;
T. Kim and C. O. Lee, ``Solutions in IR modified Ho\v{r}ava-Lifshitz gravity'', arXiv:1002.0784;
T. H. Lee, ``Global monopole asymptotic solutions in Ho\v{r}ava gravity'', Class. Quant. Grav. 27 (2010) 247001, arXiv:1003.1863;
S. S. Kim, T. Kim and Y. Kim, ``Surplus Solid Angle in Ho\v{r}ava-Lifshitz Gravity'', J. Korean Phys. Soc. 57 (2010) 634;
H. B. Kim and Y. Kim, ``Surplus angle and sign-flipped Coulomb force in projectable Ho\v{r}ava-Lifshitz gravity'', Phys. Rev. D82 (2010) 103512, arXiv:1009.1201.

\bibitem{moresolutions}
R. G. Cai, Y. Liu, and Y. W. Sun, ``On the z=4 Ho\v{r}ava-Lifshitz gravity'', JHEP 0906 (2009) 010, arXiv:0904.4104;
Y. S. Myung and Y. W. Kim, ``Thermodynamics of Ho\v{r}ava-Lifshitz black holes'', Eur. Phys. J. C68 (2010) 265, arXiv:0905.0179;
R. G. Cai, L. M. Cao and N. Ohta, ``Thermodynamics of black holes in Ho\v{r}ava-Lifshitz gravity'', Phys. Lett. B679 (2009) 504, arXiv:0905.0751;
A. Ghodsi, ``Toroidal solutions in Ho\v{r}ava gravity'', arXiv:0905.0836;
Y. S. Myung, ``Thermodynamics of black holes in the deformed Ho\v{r}ava-Lifshitz gravity'', Phys. Lett. B678 (2009) 127, arXiv:0905.0957;
A. Ghodsi and E. Hatefi,``Extremal rotating solutions in Ho\v{r}ava gravity'', Phys. Rev. D81 (2010) 044016, arXiv:0906.1237;
A. Castillo and A. Larranaga, ``Entropy for black holes in the deformed Ho\v{r}ava-Lifshitz gravity'', arXiv:0906.4380;
H. W. Lee, Y. W. Kim and Y. S. Myung, ``Extremal black holes in the Ho\v{r}ava-Lifshitz gravity'', Eur. Phys. J. C68 (2010) 255, arXiv:0907.3568;
Y. S. Myung, ``Entropy of black holes in the deformed Ho\v{r}ava-Lifshitz gravity'', Phys. Lett. B684 (2010) 158, arXiv:0908.4132;
I. Cho and G. Kang, ``Four dimensional string solutions in Ho\v{r}ava-Lifshitz gravity'', JHEP 1007 (2010) 034, arXiv:0909.3065;
R. G. Cai and N. Ohta, ``Horizon thermodynamics and gravitational field equations in Ho\v{r}ava-Lifshitz gravity'', Phys. Rev. D81 (2010) 084061, arXiv:0910.2307;
T. Harada, U. Miyamoto and N. Tsukamoto, ``Uniqueness of static spherically symmetric vacuum solutions in the IR limit of Ho\v{r}ava-Lifshitz gravity'', Int. J. Mod. Phys. D20 (2011) 111, arXiv:0911.1187;
E. Kiritsis, ``Spherically symmetric solutions in modified Ho\v{r}ava-Lifshitz gravity'', Phys. Rev. D81 (2010) 044009, arXiv:0911.3164;
J. Greenwald, A. Papazoglou and A. Wang, ``Black holes and stars in Ho\v{r}ava-Lifshitz theory with projectability condition'', Phys. Rev. D81 (2010) 084046, arXiv:0912.0011;
R. G. Cai and A. Wang, ``Singularities in Ho\v{r}ava-Lifshitz theory'', Phys. Lett. B686 (2010) 166, arXiv:1001.0155;
R. Biswas and S. Chakraborty, ``Geometry of the thermodynamics of the black holes in Ho\v{r}ava-Lifshitz gravity'', Gen. Rel. Grav. 43 (2010) 41;
S. W. Wei, Y. X. Liu, Y. Q. Wang and H. Guo, ``Thermodynamic geometry of black hole in the deformed Ho\v{r}ava-Lifshitz gravity'', arXiv:1002.1550;
Y. S. Myung, ``Lifshitz black holes in the Ho\v{r}ava-Lifshitz gravity'', Phys. Lett. B690 (2010) 534, arXiv:1002.4448;
G. Koutsoumbas, E. Papantonopoulos, P. Pasipoularides and M. Tsoukalas, ``Black hole solutions in 5D Ho\v{r}ava-Lifshitz gravity'', Phys. Rev. D81 (2010) 124014, arXiv:1004.2289;
W. Janke, D. A. Johnston and R. Kenna, ``Geometrothermodynamics of the Kehagias-Sfetsos black hole'', J. Phys. A: Math. Theor. 43 (2010) 425206, arXiv:1005.3392;
G. Koutsoumbas and P. Pasipoularides, ``Black hole solutions in Ho\v{r}ava-Lifshitz gravity with cubic terms'', Phys. Rev. D82 (2010) 044046, arXiv:1006.3199;
R. Biswas and S. Chakraborty, ``Black hole thermodynamics in Ho\v{r}ava-Lifshitz gravity and the related geometry'', Astrophys. Space Sci. 332 (2011) 193;
C. R. Arg\"{u}elles and N. E. Grandi, ``Domain Wall solutions to Ho\v{r}ava gravity'', arXiv:1008.1915;
H. W. Lee, Y. W. Kim and Y. S. Myung, ``Slowly rotating black holes in the Ho\v{r}ava-Lifshitz gravity'', Eur. Phys. J. C70 (2010) 367, arXiv:1008.2243;
A. N. Aliev and \c{C}. \c{S}ent\"{u}rk, ``Slowly rotating black hole solutions to Ho\v{r}ava-Lifshitz gravity'', Phys. Rev. D82 (2010) 104016, arXiv:1008.4848;
J. Alexandre and P. Pasipoularides, ``Spherically symmetric solutions in covariant Ho\v{r}ava-Lifshitz gravity'', Phys. Rev. D83 (2011) 084030, arXiv:1010.3634;
J. Greenwald, V. H. Satheeshkumar and A. Wang, ``Black holes, compact objects and solar system tests in non-relativistic general covariant theory of gravity'',  JCAP 12 (2010) 007, arXiv:1010.3794;
D. Capasso, ``Spherical symmetric solutions in Ho\v{r}ava-Lifshitz gravity and their properties'', Phys. Rev. D82 (2010) 124058, arXiv:1010.4326;
Q. J. Cao, Y. X. Chen and K. N. Shao, ``Black hole phase transitions in Ho\v{r}ava-Lifshitz gravity'', Phys. Rev. D83 (2011) 064015, arXiv:1010.5044;
M. Wang, S. Chen and J. Jing, ``Second-order phase transition of Kehagias-Sfetsos black hole in deformed Ho\v{r}ava-Lifshitz gravity'', Phys. Lett. B695 (2011) 401, arXiv:1012.0645;
H. Quevedo, A. Sanchez, S. Taj and A. Vazquez, ``Geometrothermodynamics in Ho\v{r}ava-Lifshitz gravity'', arXiv:1101.4494.

\bibitem{projectable}
J. Z. Tang and B. Chen, ``Static spherically symmetric solutions to modified Ho\v{r}ava-Lifshitz gravity with projectability condition'', Phys. Rev. D81 (2010) 043515, arXiv:0909.4127.

\bibitem{shift1}
E. Kiritsis and G. Kofinas, ``On Ho\v{r}ava-Lifshitz ``black holes'''', JHEP 1001 (2010) 122, arXiv:0910.5487.

\bibitem{shift2}
D. Capasso and A. P. Polychronakos, ``General static spherically symmetric solutions in Ho\v{r}ava gravity'', Phys. Rev. D81 (2010) 084009, arXiv:0911.1535.

\bibitem{chargedprojectable}
J. Z. Tang, ``Static charged black hole solutions in Ho\v{r}ava-Lifshitz gravity'', arXiv:0911.3849.

\bibitem{mass}
Y. S. Myung, ``ADM mass and quasilocal energy of black hole in the deformed Ho\v{r}ava-Lifshitz gravity'', Phys. Lett. B685 (2010) 318, arXiv:0912.3305;
M. Wang, J. Jing, C. Ding and S. Chen, ``First law of thermodynamics in IR Modified Ho\v{r}ava-Lifshitz gravity'', Phys. Rev. D81 (2010) 083006, arXiv:0912.4832.

\bibitem{ricciadm}
C. W. Misner, K. S. Thorne and J. A. Wheeler, ``Gravitation'', W. H. Freeman (1973) 519;
C. Gao, ``Modified gravity in Arnowitt-Deser-Misner formalism'', Phys. Lett. B684 (2010) 85, arXiv:0905.0310.

\bibitem{dstemp}
R. Bousso and S. W. Hawking, ``Pair creation of black holes during inflation'', Phys. Rev. D54 (1996) 6312, arXiv:gr-qc/9606052;
A. Corichi and A. Gomberoff, ``Black holes in de Sitter space: masses, energies and entropy Bounds'', Phys. Rev. D69 (2004) 064016, arXiv:hep-th/0311030.

\bibitem{wald}
R. M. Wald, ``Black hole entropy is Noether charge'', Phys. Rev. D48 (1993) R3427, arXiv:gr-qc/9307038; T. Jacobson, G. Kang and R. C. Myers, ``On black hole entropy'', Phys. Rev. D49 (1994) 6587, arXiv:gr-qc/9312023; V. Iyer and R. M. Wald, ``Some properties of Noether charge and a proposal for dynamical black hole entropy'', Phys. Rev. D50 (1994) 846, arXiv:gr-qc/9403028; T. Jacobson, G. Kang and R. C. Myers, ``Black hole entropy in higher curvature gravity'', arXiv:gr-qc/9502009.

\bibitem{mass_charge}
N. Arkani-Hamed, L. Motl, A. Nicolis and C. Vafa, ``The string landscape, black holes and gravity as the weakest force'', JHEP 0706 (2007) 060, arXiv:hep-th/0601001;
Y. Kats, L. Motl and M. Padi, ``Higher-order corrections to mass-charge relation of extremal black holes'', JHEP 0712 (2007) 068, arXiv:hep-th/0606100.

\bibitem{rnds}
L. J. Romans, ``Supersymmetric, cold and lukewarm black holes in cosmological Einstein-Maxwell theory'', Nucl. Phys. B383 (1992) 395, arXiv:hep-th/9203018;
R. B. Mann and S. F. Ross, ``Cosmological production of charged black hole pairs'', Phys. Rev. D52 (1995) 2254, arXiv:gr-qc/9504015;
R. B. Mann, ``Charged topological black hole pair creation'', Nucl. Phys. B516 (1998) 357, arXiv:hep-th/9705223;
J. H. Cho and S. Nam, ``Non-supersymmetric attractor with the cosmological constant'', JHEP 0707 (2007) 011, arXiv:0705.2892.

\bibitem{tmg}
S. Deser, R. Jackiw and S. Templeton, ``Topologically massive gauge theories'', Ann. Phys. 140 (1982) 372; erratum: ibid. 185 (1998), 406;
S. Deser, R. Jackiw and S. Templeton, ``Three-dimensional massive gauge theories'', Phys. Rev. Lett. 48 (1982) 975.

\bibitem{csmecs}
S. Deser and R. Jackiw, ``Higher derivative Chern-Simons Extensions'', Phys. Lett. B451 (1999) 73, arXiv:hep-th/9901125.

\end{thebibliography}

\documentclass[12pt]{article}
\usepackage{amssymb,amsmath}

\newcommand{\half} {\frac{1}{2}}
\newcommand{\quart} {\frac{1}{4}}
\newcommand{\abs} [1] {\left|#1\right|}
\newcommand{\dash} {\nobreakdash-\hspace{0pt}}
\renewcommand{\eqref} [1] {(\ref{#1})}
\newcommand{\LW} {\Lambda_W}

\begin{document}

\begin{titlepage}
\hfill
\vbox{
    \halign{#\hfil         \cr
           } 
      }  
\vspace*{20mm}

\begin{center}
{\Large {\bf Black Holes in Ho\v{r}ava Gravity with\\ Higher Derivative Magnetic Terms}\\} \vspace*{15mm}

{\sc Eyal Gruss} \footnote{e-mail: {\tt eyalgruss@gmail.com}}

\vspace*{1cm}
{\it Raymond and Beverly Sackler School of Physics and Astronomy,\\
Tel-Aviv University, Tel-Aviv 69978, Israel.\\}

\end{center}

\vspace*{8mm}

\begin{abstract}
We consider Ho\v{r}ava gravity coupled to Maxwell and higher derivative magnetic terms. We construct static spherically symmetric black hole solutions in the low-energy approximation. We calculate the horizon locations and temperatures in the near-extremal limit, for asymptotically flat and (anti\dash)de Sitter spaces. We also construct a detailed balanced version of the theory, for which we find projectable and non-projectable, non-perturbative solutions.
\end{abstract}
\vskip 0.8cm

May 2010.

\end{titlepage}

\setcounter{footnote}{0}

\section{Introduction}

Ho\v{r}ava has recently proposed a theory of gravity which is power-counting renormalizable and possibly unitary \cite{horava}. This is achieved by assigning different scaling dimensions to time and space, at the cost of breaking Lorentz and diffeomorphism invariance at high energies. In four-dimensional Ho\v{r}ava gravity, one usually considers an anisotropic scaling exponent of $z=3$, such that $t\sim x^z$. This means that the Lagrangian should include higher spatial derivative terms up to sixth order. One may also consider couplings to scalar \cite{scalarvector1,scalar1,scalar2,scalarvector2} and vector \cite{scalarvector1,scalarvector2,magnetic,vector} fields with higher derivatives terms. There is an ongoing discussion regarding possible flaws of the theory. We refer to the latest reviews on these subjects \cite{problems}, and references therein.

The study of Black hole solutions and thermodynamics in general relativity with higher curvature corrections has been pioneered by \cite{CMP}. There, large black hole solutions were obtained by treating the higher curvature terms as a small perturbation. Black hole solutions in Ho\v{r}ava gravity and their thermodynamics have been discussed in \cite{deformed,charged,moresolutions,projectable,shift1,shift2,chargedprojectable,mass}. In particular, \cite{charged,chargedprojectable} discusses charged black holes in Ho\v{r}ava gravity coupled to the usual Maxwell term. In this work we add higher derivative magnetic terms to the mix, and discuss black hole solutions and thermodynamics.

The paper is organized as follows: In section $2$, we discuss Ho\v{r}ava-Maxwell theory with higher derivative magnetic terms. In section $3$, we discuss static spherically symmetric black hole solutions in the low-energy approximation. In section $3.1$, we calculate the horizon locations and temperatures for near-extremal asymptotically flat black holes. In section $3.2$, we do the same for near-cold asymptotically (anti\dash)de Sitter black holes. In section $4$, we discuss a detailed balanced version of the theory, and find non-perturbative solutions.

\section{Ho\v{r}ava-Maxwell Theory}

We use the ADM formalism where the metric is parameterized as
\begin{equation}
ds^2=-N^2dt^2+g_{ij}(dx^i+N^idt)(dx^j+N^jdt) \ ,
\end{equation}
where $i,j$ run over spatial coordinates. The extrinsic curvature is given by
\begin{equation}
K_{ij}=\frac{1}{2N}(\dot{g}_{ij}-\nabla_iN_j-\nabla_jN_i) \ , \qquad K\equiv K^i_{\phantom{i}i} \ ,
\end{equation}
where dot denotes a derivative with respect to $t$.
The four-dimensional Ricci scalar may be decomposed (see e.g. \cite{ricciadm}) as
\begin{equation}
\label{R4}
{}^{(4)}R=R+K_{ij}K^{ij}-K^2+(\textrm{covariant derivative}) \ ,
\end{equation}
where $R$ is the three-dimensional Ricci scalar. The last term contributes a total derivative in the Einstein-Hilbert Lagrangian, and may be dropped in this case. Note however, that this will not be true in general for higher curvature terms. Thus, Ho\v{r}ava gravity, which is of higher curvature in $R$, will differ from the more common ${}^{(4)}R$ higher curvature theories, even for static solutions. The Cotton tensor is given by
\begin{equation}
\label{cotton}
C_{ij}=\frac{\epsilon^{ikl}}{\sqrt{g}}\nabla_k\left(R^j_{\phantom{j}l}-\quart R\delta^j_l\right) \ ,
\end{equation}
where $\epsilon^{123}=1$, $g\equiv\mathrm{det}(g_{ij})$, and $R_{ij}$ is the three-dimensional Ricci tensor.

Ho\v{r}ava's ``detailed balance'' Lagrangian density \cite{horava} is given by: $\mathcal{L}_{Ho\check{r}ava}=\mathcal{L}_0+\mathcal{L}_1$,
with
\begin{eqnarray}
\label{Lhorava}
\frac{1}{N\sqrt{g}}\mathcal{L}_0&=&\frac{2}{\kappa^2}(K_{ij}K^{ij}-\lambda K^2)+\frac{\kappa^2\mu^2}{8(1-3\lambda)}(\LW R-3\LW^2) \ , \nonumber\\*
\frac{1}{N\sqrt{g}}\mathcal{L}_1&=&\frac{\kappa^2\mu^2(1-4\lambda)}{32(1-3\lambda)}R^2-\frac{\kappa^2\mu^2}{8}R_{ij}R^{ij}+\frac{\kappa^2\mu}{2w^2}\frac{\epsilon^{ijk}}{\sqrt{g}}R_{il}\nabla_jR^l_{\phantom{l}k}-\frac{\kappa^2}{2w^4}C_{ij}C^{ij} \ ,
\end{eqnarray}
where $\kappa$, $\lambda$, $\mu$, $\LW$ and $w$ are constant parameters. Comparing $\mathcal{L}_0$ to the Einstein-Hilbert Lagrangian, these parameters are related to the speed of light, Newton's constant and the cosmological constant by
\begin{equation}
c=\frac{\kappa^2\mu}{4}\sqrt{\frac{\LW}{1-3\lambda}} \ , \qquad G_N=\frac{\kappa^2}{32\pi c} \ , \qquad \Lambda=\frac{3}{2}\LW \ .
\end{equation}
We want this theory to reduce to general relativity in the low-energy limit. We therefore first set $\lambda=1$. Choosing units so that $c=1$, shows the coefficients of the $R^2$, $R_{ij}R^{ij}$ terms in $\mathcal{L}_1$ to be inversely proportional to ${\LW}$, with no additional free parameters. In order to allow both the cosmological constant term and the higher curvature terms to be small, we must deform the theory to break the relation between these terms \cite{horava,deformed}. As long as we do not change the form of the six-derivative $C_{ij}C^{ij}$ term, this is considered a soft violation of the detailed balance condition. We will write this as: $\mathcal{L}_{deformed}=\mathcal{L}_0+\epsilon a_0\LW\mathcal{L}_1$, where $a_0$ is a constant parameter, and $\epsilon>0$ will be used later as a small expansion parameter. In this parameterization $\LW$ factors out, leaving the now independent usual cosmological constant term (see e.g. \eqref{Lreduced}).

We now consider electromagnetic interactions. Maxwell's Lagrangian density in the ADM decomposition takes the form:
\begin{equation}
\frac{\gamma}{N\sqrt{g}}\mathcal{L}_2=\frac{2}{N^2}g^{ij}(F_{ti}-N^kF_{ki})(F_{tj}-N^lF_{lj})-F_{ij}F^{ij} \ ,
\end{equation}
where $\gamma=16\pi Gc^{-1}$ fixes the normalization in relation to the gravity part, and with the field strength given by
\begin{equation}
F_{ti}=\dot{A}_i-\partial_iA_t \ , \qquad F_{ij}=\partial_iA_j-\partial_jA_i \ ,
\end{equation}
where $(A_t,A_i)$ is the vector potential. We will add minimally coupled higher derivative magnetic terms as considered in \cite{scalarvector2,magnetic}:
\begin{equation}
\label{L3}
\frac{\gamma}{N\sqrt{g}}\mathcal{L}_3=-a_1\nabla_kF_{ij}\nabla^kF^{ij}-a_2\nabla_k\nabla_lF_{ij}\nabla^k\nabla^lF^{ij} \ ,
\end{equation}
where $a_1$, $a_2$ are constant parameters. \cite{magnetic} discusses constraints on the magnitude of $a_2$. Note that contrary to the case of \cite{scalarvector2,magnetic} which consider cosmological backgrounds, terms related by integration by parts to the RHS of \eqref{L3} are relevant since $N$ is space-dependant (in the non-projectable case). Nevertheless such terms are not considered here. Note also that in the spherically symmetric case considered later, $\nabla_jF^{ij}=0$ and $\triangle F^{ij}=\frac{1}{2g}\epsilon^{ijk}\epsilon^{lmn}R_{kl}F_{mn}$. Therefore, we have not considered here terms containing such derivatives. In section $4$, we consider a detailed balanced version of Ho\v{r}ava-Maxwell theory, which does include such terms.

We take our overall Lagrangian to be
\begin{equation}
\mathcal{L}=\mathcal{L}_0+\epsilon a_0\LW\mathcal{L}_1+\mathcal{L}_2+\epsilon\mathcal{L}_3 \ .
\end{equation}
Finally, we make the large black hole (or low-energy) approximation, in the spirit of \cite{CMP}. We require that the gravitational and magnetic higher derivative terms are small compared to $R$ and $F_{ij}F^{ij}$, respectively. In our analysis this will usually mean
\begin{equation}
\epsilon\abs{a_0}\ll r_{hor}^2 \ , \qquad \epsilon\abs{a_1}p^2\ll r_{hor}^4 \ , \qquad \epsilon\abs{a_2}p^2\ll r_{hor}^6 \ ,
\end{equation}
where $r_{hor}$ is the location of the horizon of interest, and $p$ is the magnetic charge. However, for asymptotically anti-de Sitter solutions with a large (negative) cosmological constant: $\abs{\LW}r_{hor}^2\gg1$, the conditions instead read
\begin{equation}
\epsilon\abs{a_0\LW}\ll1 \ , \qquad \epsilon\abs{a_1}p^2\ll r_{hor}^4 \ , \qquad \epsilon\abs{a_2\LW}p^2\ll r_{hor}^4 \ .
\end{equation}
This may be seen from \eqref{phi}. We will construct perturbative solutions of this theory to first order in $\epsilon$. Our calculations were done using Maple with GRTensor.

\section{Static Spherically Symmetric Black Holes}

We consider static spherically symmetric solutions, where the metric takes the form:
\begin{equation}
\label{gSSS}
ds^2=\Big(-N(r)^2+N_r(r)^2f(r)\Big)dt^2+2N_r(r)dtdr+\frac{1}{f(r)}dr^2+r^2(d\theta^2+sin^2\theta d\phi^2) \ ,
\end{equation}
where we have retained the shift function $N_r$ as discussed in \cite{projectable,shift1,shift2}. The vector potential takes the form:
\begin{equation}
\label{ASSS}
A_t=A(r) \ , \qquad A_r=A_\theta=0 \ , \qquad A_\phi=-p\cos\theta \ .
\end{equation}
Substituting the above ansatz into the Lagrangian density we get
\begin{eqnarray}
\label{Lreduced}
\frac{\sqrt{f}}{N}r^6\mathcal{L}_{reduced}&\propto&3\LW r^8+2f'r^7+2fr^6-2r^6+\epsilon a_0(2f'fr^5-2f'r^5-f^2r^4+2fr^4-r^4)+\nonumber\\*
&&{}-\frac{2f}{N^2}A'^2r^8+2p^2r^4+12\epsilon a_1p^2fr^2+3\epsilon a_2p^2(f'^2r^2-12f'fr+60f^2)+\nonumber\\*
&&{}+\frac{2f}{c^2N^2}(2N_r'N_rfr^7+N_r^2f'r^7+N_r^2fr^6) \ ,
\end{eqnarray}
where prime denotes a derivative with respect to $r$, and we have omitted an overall factor independent or $r$. Varying $\mathcal{L}_{reduced}$ with respect to $N$, $N_r$, $f$ and $A$, yields the equations of motion, the latter giving
\begin{equation}
A'=\frac{Nq}{\sqrt{f}r^2} \ ,
\end{equation}
where $q$ is the electric charge.

The variation by $N_r$ gives
\begin{equation}
\label{EOMNr}
\frac{N_rf^{3/2}r}{N}\left(\frac{2N'}{N}-\frac{f'}{f}\right)=0 \ .
\end{equation}
Let us first consider the case $N_r\neq0$. The motivation here is to obtain solutions to the projectable version of Ho\v{r}ava gravity, where $N=N(t)$ \cite{horava,projectable,chargedprojectable}. In our static case, $N$ must then be constant, implying that also $f$ is constant due to \eqref{EOMNr}. For $a_1=a_2=0$, uncharged and charged solutions of $N_r$ were discussed in \cite{projectable} and \cite{chargedprojectable}, respectively.\footnote{See e.g. \eqref{NrDB} with $\beta=0$, assuming $\epsilon a_0\LW=1$.} For general $a_1\neq0$ or $a_2\neq0$, we find there are no magnetically charged solutions to the equations of motion. \cite{shift2} discusses black hole thermodynamics when $N_r\neq0$. From here on we assume $N_r=0$.

The zeroth order solution is the Reissner-Nordstr\"om-(anti\dash)de Sitter black hole:
\begin{equation}
\label{N02}
N_0^2=f_0=-\frac{\LW}{2}r^2+1-\frac{2m_0}{r}+\frac{q^2+p^2}{r^2} \ ,
\end{equation}
where $m_0$ is the mass at zeroth order (see however \cite{mass}). For $a_1=a_2=0$, uncharged and charged non-perturbative solutions were obtained in \cite{deformed} and \cite{charged,chargedprojectable}, respectively.\footnote{See e.g. \eqref{N2DB} with $\beta=0$, assuming $\epsilon a_0\LW=1$.} For the general case, we will look for solutions to first order in $\epsilon$, in the form:
\begin{equation}
f=f_0(1+\epsilon\varphi) \ , \qquad N^2=f(1+\epsilon\eta) \ .
\end{equation}
Note that in the expression for $N^2$ we use $f$ rather than $f_0$, and truncate above order $\epsilon$. The solution reads
\begin{eqnarray}
\label{phi}
\varphi&=&\Bigg[105a_0\Big(\LW^2r^{12}-4\LW(q^2+p^2)r^8+16m_0^2r^6-16m_0(q^2+p^2)r^5+4(q^2+p^2)^2r^4\Big)+\nonumber\\*
&&{}+168a_1p^2\Big(15\LW r^8-10r^6+15m_0r^5-6(q^2+p^2)r^4\Big)+\nonumber\\*
&&{}+8a_2p^2\Big(-1575\LW^2r^8+2520\LW r^6-4095\LW m_0r^5+1764\LW(q^2+p^2)r^4-1890r^4+\nonumber\\*
&&{}\phantom{+8a_2p^2\Big(}+6930m_0r^3-6570m_0^2r^2-3240(q^2+p^2)r^2+6300m_0(q^2+p^2)r+\nonumber\\*
&&{}\phantom{+8a_2p^2\Big(}-1540(q^2+p^2)^2\Big)+\nonumber\\*
&&{}+1680m_0\delta r^9\Bigg]\Bigg/\Bigg[420\Big(\LW r^{12}-2r^{10}+4m_0r^9-2(q^2+p^2)r^8\Big)\Bigg] \ ,
\end{eqnarray}
and
\begin{equation}
\eta=3p^2\left(\frac{-a_1+9a_2\LW}{r^4}-\frac{10a_2}{r^6}+\frac{96a_2m_0}{7r^7}-\frac{4a_2(q^2+p^2)}{r^8}\right) \ ,
\end{equation}
where $\delta$ is an integration constant, and we have assumed the $\epsilon$-corrections vanish at infinity. Since in general $\eta\neq0$, $N^2$ is not proportional to $f$. Thus there is no transformation of coordinates that would yield a corresponding projectable solution \cite{projectable}, as indeed expected in light of the above discussion.

Looking at $N(r)^2$, one sees that the $\delta$-term is just an explicit $\epsilon$-correction to the mass: $m=m_0(1+\epsilon\delta)$. We assume the horizon locations of the black hole are given by the roots of $g_{tt}=N(r)^2=0$, as in the relativistic case (see \cite{shift1} for a discussion of the notion of horizon in Ho\v{r}ava gravity). One may set $\delta$ to make $\varphi(r)$ and $\eta(r)$ nonsingular on the unperturbated black hole event horizon, as done in \cite{CMP}. This would mean the event horizon location does not get an $\epsilon$-correction. However, the mass then has a fixed $\epsilon$-correction, and this depends nontrivially on the zeroth order mass $m_0$. Also, other horizons of the solution do get $\epsilon$-corrections. We will take the alternate route of leaving $\delta$ free, and finding all horizon locations perturbatively, designating: $r_{hor}=r_0(1+\epsilon\rho)$, with $\epsilon\abs{\rho}\ll1$. Note that in the near-horizon region, $f(r)$ and $N^2(r)$ now get an additional $\epsilon$-correction due to $\rho$ in $f_0(r)$, and it is the overall correction that should be small.

The Hawking temperature of our static spherically symmetric black hole (with $N_r=0$, otherwise see \cite{shift2}) is given by
\begin{equation}
T=\frac{\abs{g_{tt}'}}{4\pi\sqrt{-g_{tt}g_{rr}}}\Bigg|_{r=r_{hor}}=\frac{\abs{N'}\sqrt{\abs{f}}}{2\pi}\Bigg|_{r=r_{hor}}
\ .
\end{equation}
Note that for asymptotically de Sitter black holes, the temperature may need to include the Bousso-Hawking normalization factor \cite{dstemp}, not considered here. We do not calculate the Bekenstein-Hawking area law entropy, as it is expected that the entropy would be corrected by the higher derivative terms. In fully diffeomorphism invariant theories, this is computed by the Wald entropy \cite{wald}. However, as mentioned also in \cite{shift1}, it is not clear how to do this in our diffeomorphism breaking theory. In the following we will analyze the horizon locations and temperatures for some simplified cases of black holes.

\subsection{Asymptotically Flat Black Holes}
We first consider asymptotically flat black holes with $\LW=0$ (holding $c$ finite). The zeroth order inner and outer horizons are given by
\begin{equation}
r_0^{in}=m_0-\sqrt{m_0^2-q^2-p^2} \ , \qquad r_0^{out}=m_0+\sqrt{m_0^2-q^2-p^2} \ .
\end{equation}
The expressions for the first order corrections are lengthy, and to be concise we will consider only the extremal and near-extremal cases. The extremal limit is determined by: $r^{in}=r^{out}$, which is satisfied at zeroth order by
\begin{equation}
r_0^{extr}=m_0^{extr}=\sqrt{q^2+p^2} \ .
\end{equation}
At this extremum point, the first order correction to the horizon location will not affect $N(r)^2$ to first order in $\epsilon$. In order to satisfy $N(r)^2=0$, one must set
\begin{equation}
\delta^{extr}=\frac{-105a_0(q^2+p^2)^2+42a_1(q^2+p^2)p^2+20a_2p^2}{420(q^2+p^2)^3} \ ,
\end{equation}
correcting the extremal mass. It is conjectured that the mass-charge ratio of extremal black holes is decreased by higher curvature corrections \cite{mass_charge}, implying: $\delta^{extr}<0$. The first order correction to the extremal horizon may then be got either by taking the limit of the non-extremal case, or by satisfying: $\left(N(r)^2\right)'=0$, that is, having the temperature vanish. It reads
\begin{equation}
\rho^{extr}=-\frac{a_0}{4(q^2+p^2)} \ .
\end{equation}
Note that the extremal horizon does not depend on $a_1$, $a_2$.

In the near-extremal approximation we parameterize the mass as
\begin{equation}
m=m^{extr}\sqrt{1+\Delta^2} \ , \qquad m^{extr}=m_0^{extr}(1+\epsilon\delta^{extr}) \ ,
\end{equation}
and expand all expressions to leading order in $\Delta\ll1$. The horizon locations satisfying: $N(r)^2=0+O(\Delta^3)$, are given at zeroth order in $\epsilon$ by
\begin{equation}
r_0^{in}=\sqrt{q^2+p^2}(1-\Delta)+O(\Delta^2) \ , \qquad r_0^{out}=\sqrt{q^2+p^2}
(1+\Delta)+O(\Delta^2) \ ,
\end{equation}
and at first order in $\epsilon$ by
\begin{eqnarray}
\rho^{in}&=&\frac{-105a_0(q^2+p^2)^2-336a_1(q^2+p^2)p^2\Delta-100a_2p^2\Delta}{420(q^2+p^2)^3}+O(\Delta^2) \ , \nonumber\\ \rho^{out}&=&\frac{-105a_0(q^2+p^2)^2+336a_1(q^2+p^2)p^2\Delta+100a_2p^2\Delta}{420(q^2+p^2)^3}+O(\Delta^2) \ .
\end{eqnarray}
The Hawking temperature is
\begin{equation}
T^{out}=\frac{\Delta}{2\pi\sqrt{q^2+p^2}}\left(1+\epsilon\frac{
525a_0(q^2+p^2)^2-294a_1(q^2+p^2)p^2-80a_2p^2}{420(q^2+p^2)^3}\right)+O(\Delta^2+\epsilon^2) \ .
\end{equation}

\subsection{Asymptotically (Anti\dash)de Sitter Black Holes}
We now consider asymptotically (anti\dash)de Sitter black holes with $\LW\neq0$. In the de Sitter case, having $\LW>0$ (holding $c$ positive by analytic continuation), there exists a cosmological horizon denoted: $r^{cosm}$. Let us recall the classification of the limiting cases of Reissner-Nordstr\"om-de Sitter black holes \cite{rnds,dstemp}:
\begin{enumerate}
\item[(a)] lukewarm: $r^{in}<r^{out}<r^{cosm}$ and $T^{out}=T^{cosm}$.
\item[(b)] charged Nariai: $r^{in}<r^{out}=r^{cosm}$.
\item[(c)] cold: $r^{in}=r^{out}<r^{cosm}$.
\item[(d)] ultracold: $r^{in}=r^{out}=r^{cosm}$.
\end{enumerate}
In the anti-de Sitter case, $\LW<0$, and the only relevant limit is the cold black hole: $r^{in}=r^{out}$. Together with the flat case, these limiting cases (and their ``near'' versions) have a concise zeroth order form for the horizon location. We will first consider the cold and ultracold limits. We will then extend to the near-cold black hole: $r^{in}\lesssim r^{out}$. For the latter to be concise also at first order in $\epsilon$, we will consider the near-flat approximation: $\abs{\LW}(q^2+p^2)\ll1$.

The zeroth order cold black hole metric function is of fourth degree in $r$, and has a double zero. It may be written as
\begin{equation}
\label{doublezero}
\left(N_0^{cold}\right)^2=-\frac{\LW}{2r^2}(r-A)(r-B)^2(r-C) \ ,
\end{equation}
where $A$, $B$, $C$ are constants. Equating \eqref{doublezero} with \eqref{N02} and comparing powers of $r$, one gets
\begin{eqnarray}
A+2B+C&=&0 \ , \nonumber\\
3\LW B^2+2\LW BC+\LW C^2-2&=&0 \ , \nonumber\\
\LW B^3-B+m_0&=&0 \ , \nonumber\\
3\LW B^4-2B^2+2(q^2+p^2)&=&0 \ .
\end{eqnarray}
This gives
\begin{equation}
\label{r0cold}
r_0^{cold}=B=\sqrt{\frac{2(q^2+p^2)}{\sqrt{1-6\LW(q^2+p^2)}+1}} \ ,
\end{equation}
and
\begin{equation}
m_0^{cold}=\frac{B}{3}\left(2+\sqrt{1-6\LW(q^2+p^2)}\right) \ ,
\end{equation}
where for $\LW>0$ we have assumed $B\leq C$, and in that case we have also
\begin{equation}
r_0^{cosm}=C=-B+\sqrt{\frac{4+2\sqrt{1-6\LW(q^2+p^2)}}{3\LW}} \ .
\end{equation}
As in the extremal flat case, to satisfy $N(r)^2=0$ to first order in $\epsilon$, one must correct the cold black hole mass by setting
\begin{eqnarray}
\delta^{cold}&=&\Bigg[70a_0(q^2+p^2)^2\Big(-7+12\LW(q^2+p^2)-2\sqrt{1-6\LW(q^2+p^2)}\Big)+\nonumber\\*
&&{}+126a_1(q^2+p^2)p^2\Big(1-18\LW(q^2+p^2)+\sqrt{1-6\LW(q^2+p^2)}\Big)+\nonumber\\*
&&{}+3a_2p^2\Big(20-318\LW(q^2+p^2)+2844\LW^2(q^2+p^2)^2-23\sqrt{1-6\LW(q^2+p^2)}+\nonumber\\*
&&{}\phantom{+3a_2p^2\Big(}+43\left(1-6\LW(q^2+p^2)\right)^{3/2}\Big)\Bigg]\frac{1+\sqrt{1-6\LW(q^2+p^2)}}{1680(q^2+p^2)^3\left(2+\sqrt{1-6\LW(q^2+p^2)}\right)} \ , \nonumber\\*
\end{eqnarray}
The first order correction to the cold horizon location reads
\begin{equation}
\label{rhocold}
\rho^{cold}=-a_0\frac{1+\sqrt{1-6\LW(q^2+p^2)}}{8(q^2+p^2)\sqrt{(1-6\LW(q^2+p^2)}} \ .
\end{equation}
Here again the cold horizon does not depend on $a_1$, $a_2$. The full expressions for the cosmological horizon location and temperature are more cumbersome, and will be given later in the near-flat limit.

The ultracold limit is determined by: $r^{cold}=r^{cosm}$, which is satisfied at zeroth order by
\begin{equation}
\label{UC}
\Lambda_0^{UC}=\frac{1}{6(q^2+p^2)} \ ,
\end{equation}
and
\begin{equation}
r_0^{UC}=\frac{3}{2}m_0^{UC}=\sqrt{2(q^2+p^2)} \ ,
\end{equation}
where we have denoted: $\LW=\Lambda_0(1+\epsilon\upsilon)$, and we expand in $\epsilon$ before taking the ultracold limit. Note that the first order corrections to the mass and horizon location, now also get a contribution from their respective zeroth order terms due to $\upsilon$. Thus the $\epsilon$-correction in the expansion of \eqref{r0cold}, serves to cancel out the divergency of \eqref{rhocold} in the limit \eqref{UC}. The overall first order terms are given by
\begin{eqnarray}
\upsilon^{UC}&=&\frac{a_0}{2(q^2+p^2)} \ , \nonumber\\
\delta^{UC}&=&\frac{-280a_0(q^2+p^2)^2-126a_1(q^2+p^2)p^2+69a_2p^2}{1680(q^2+p^2)^3} \ , \nonumber\\
\rho^{UC}&=&-\frac{a_0}{4(q^2+p^2)} \ .
\end{eqnarray}

In the near-cold approximation we parameterize the mass as
\begin{equation}
m=m^{cold}\sqrt{1+\Delta^2} \ , \qquad m^{cold}=m_0^{cold}(1+\epsilon\delta^{cold}) \ ,
\end{equation}
and expand all expressions to leading order in $\Delta\ll1$. The horizon locations satisfying: $N(r)^2=0+O(\Delta^3)$, are given at zeroth order in $\epsilon$ by
\begin{equation}
r_0^{in}=b(1-\alpha_b\Delta)+O(\Delta^2) \ , \qquad r_0^{out}=b
(1+\alpha_b\Delta)+O(\Delta^2) \ , \qquad r_0^{cosm}=c
(1+\alpha_c\Delta^2)+O(\Delta^3) \ ,
\end{equation}
where
\begin{eqnarray}
\alpha_b&=&\sqrt{\frac{1}{3}+\frac{2}{3\sqrt{1-6\LW(q^2+p^2)}}} \ , \nonumber\\
\alpha_c&=&\frac{(2+S)\left(1-S-\sqrt{4-2S-2S^2}\right)}{6\left(8-2S^2-4\sqrt{4-2S-2S^2}-S\sqrt{4-2S-2S^2}\right)} \ , \qquad S=\sqrt{1-6\LW(q^2+p^2)}\ . \nonumber\\*
\end{eqnarray}

We now take the near-flat approximation by expanding to leading order in $L=\abs{\LW}(q^2+p^2)\ll1$. The mass, the horizon locations and the temperatures of the near-cold near-flat black hole are given by
\begin{eqnarray}
m_0^{cold}&=&\sqrt{q^2+p^2}\left(1-\frac{L}{4}\right)+O(L^2) \ , \nonumber\\
r_0^{in}&=&\sqrt{q^2+p^2}\left(1+\frac{3L}{4}-\Delta-\frac{7L\Delta}{4}\right)+O(L^2+\Delta^2) \ , \nonumber\\
r_0^{out}&=&\sqrt{q^2+p^2}\left(1+\frac{3L}{4}+\Delta+\frac{7L\Delta}{4}\right)+O(L^2+\Delta^2) \ , \nonumber\\
r_0^{cosm}&=&\sqrt{q^2+p^2}\left(\sqrt{2}L^{-1/2}-1-\frac{\Delta^2}{2}\right)+O(L^{1/2}+\Delta^4) \ , \nonumber\\
\delta^{cold}&=&\delta^{extr}_{L=0}+\frac{525a_0(q^2+p^2)^2-924a_1(q^2+p^2)p^2-656a_2p^2}{840(q^2+p^2)^3}L+O(L^2) \ , \nonumber\\
\rho^{in}&=&\rho^{in}_{L=0}+\frac{105a_0(q^2+p^2)^2(-3+4\Delta)+1680a_1(q^2+p^2)p^2\Delta+1748a_2p^2\Delta}{840(q^2+p^2)^3}L+O(L^2+\Delta^2) \ , \nonumber\\
\rho^{out}&=&\rho^{out}_{L=0}-\frac{105a_0(q^2+p^2)^2(3+4\Delta)+1680a_1(q^2+p^2)p^2\Delta+1748a_2p^2\Delta}{840(q^2+p^2)^3}L+O(L^2+\Delta^2) \ , \nonumber\\
\rho^{cosm}&=&-\frac{\delta^{extr}_{L=0}L^{1/2}}{\sqrt{2}}\left(1+\frac{\Delta^2}{2}\right)+O(L+\Delta^4) \ , \nonumber\\
T^{out}&=&T^{out}_{L=0}-\frac{11L\Delta}{8\pi\sqrt{q^2+p^2}}\left(1+\epsilon\frac{5985a_0(q^2+p^2)^2-6090a_1(q^2+p^2)p^2-3464a_2p^2}{4620(q^2+p^2)^3}\right)+\nonumber\\
&&{}+O(L^2+\Delta^2+\epsilon^2) \ , \nonumber\\
T^{cosm}&=&\frac{\sqrt{2}L^{1/2}-2L-L\Delta^2}{4\pi\sqrt{q^2+p^2}}\left(1-\sqrt{2}\epsilon\delta^{extr}_{L=0}L^{1/2}\left(1+\frac{\Delta^2}{2}\right)\right)+O(L^{3/2}+\Delta^4+\epsilon^2) \ .
\end{eqnarray}
where $L=0$ designates the near-extremal flat case results of the previous section. For $\Delta=0$, we get the cold near-flat limit.

\section{A Detailed Balanced Ho\v{r}ava-Maxwell Theory}

Detailed balanced Ho\v{r}ava gravity was presented in \cite{horava}. A detailed balanced version of the theory coupled to a scalar field, was derived in \cite{scalar2}. Here we discuss a detailed balanced version of Ho\v{r}ava gravity coupled to Maxwell and higher derivative magnetic terms. Detailed balance is a symmetry condition which serves to reduce the number of terms in the Lagrangian. This allows in our case to yield analytic non-perturbative solutions.

The detailed balance Lagrangian density will be given by: $\mathcal{L}=\mathcal{L}_K-\mathcal{L}_V$. The kinetic part is the same as before:
\begin{equation}
\frac{1}{N\sqrt{g}}\mathcal{L}_K=\frac{2}{\kappa^2}(K_{ij}K^{ij}-\lambda K^2)+\frac{2}{\gamma N^2}g^{ij}(F_{ti}-N^kF_{ki})(F_{tj}-N^lF_{lj}) \ .
\end{equation}
The potential part is defined to be of the form:
\begin{equation}
\frac{1}{N\sqrt{g}}\mathcal{L}_V=\frac{\kappa^2}{8}\mathcal{G}_{ijkl}E^{ij}E^{kl}+\frac{1}{2\gamma}E_iE^i \ ,
\end{equation}
with the generalized de Witt metric:
\begin{equation}
\mathcal{G}_{ijkl}=\half(g_{ik}g_{jl}+g_{il}g_{jk})+\frac{\lambda}{1-3\lambda}g_{ij}g_{kl} \ ,
\end{equation}
and where
\begin{equation}
E^{ij}=\frac{1}{\sqrt{g}}\frac{\delta W}{\delta g_{ij}} \ , \qquad E^i=\frac{1}{\sqrt{g}}\frac{\delta W}{\delta A_i} \ ,
\end{equation}
for some three-dimensional Euclidean Lagrangian density $W$. For $z=3$, $W$ should include up to three spatial derivatives.

The gravitational part of $W$, due to \cite{horava}, is given by
\begin{equation}
\frac{1}{\sqrt{g}}W_{Ho\check{r}ava}=-\frac{1}{w^2}\frac{\epsilon^{ijk}}{\sqrt{g}}\Gamma^{m}_{\phantom{m}il}\left(\partial_j\Gamma^l_{\phantom{l}km}+\frac{2}{3}\Gamma^l_{\phantom{l}jn}
\Gamma^{n}_{\phantom{n}km}\right)+\mu(R-2\LW) \ ,
\end{equation}
where $\Gamma^{i}_{\phantom{i}jk}$ are the Christoffel symbols. This is the Lagrangian density of topological massive gravity \cite{tmg}, with the first term being the gravitational Chern-Simons term\footnote{Note that to be consistent with \eqref{Eijhor} and \eqref{cotton}, we find this term to have an opposite sign compared to \cite{horava}.}. Varying with respect to $g_{ij}$ yields
\begin{equation}
\label{Eijhor}
E^{ij}_{Ho\check{r}ava}=\frac{1}{\sqrt{g}}\frac{\delta W_{Ho\check{r}ava}}{\delta g_{ij}}=\frac{2}{w^2}C^{ij}-\mu\left(R^{ij}-\half Rg^{ij}+\LW g^{ij}\right) \ .
\end{equation}
Using this, one gets the potential part in $\mathcal{L}_{Ho\check{r}ava}$ \eqref{Lhorava}.

For the electromagnetic part of $W$, we take
\begin{equation}
\frac{1}{\sqrt{g}}W_{EM}=\frac{\epsilon^{ijk}}{\sqrt{g}}A_i\partial_jA_k+b_1F_{ij}F^{ij}+b_2\frac{\epsilon^{ijk}}{\sqrt{g}}(\triangle A_i-R^l_{\phantom{l}i}A_l)\partial_jA_k \ ,
\end{equation}
where $b_1$, $b_2$ are constant parameters. This is the three-dimensional Chern-Simons-Maxwell-extended-Chern-Simons Lagrangian density in curved space \cite{csmecs}. Varying with respect to $g_{ij}$ yields
\begin{eqnarray}
E^{ij}_{EM}=\frac{1}{\sqrt{g}}\frac{\delta W_{EM}}{\delta g_{ij}}&=&b_1\left(\half F_{kl}F^{kl}g^{ij}-2g_{kl}F^{ki}F^{lj}\right)+\nonumber\\* &&{}+b_2\frac{\epsilon^{ikl}}{\sqrt{g}}(F^{jm}\nabla_kF_{lm}+F_{lm}\nabla_kF^{jm}-g_{kn}F_{lm}\nabla^jF^{mn}) \ ,
\end{eqnarray}
where in the spherically symmetric case, the $b_2$ term vanishes. Varying with respect to the $A_i$ yields
\begin{equation}
E^i=\frac{1}{\sqrt{g}}\frac{\delta W_{EM}}{\delta A_i}=\frac{\epsilon^{ijk}}{\sqrt{g}}F_{jk}+4b_1\nabla_jF^{ij}+b_2\frac{\epsilon^{jkl}}{\sqrt{g}}\left(\delta^i_j\triangle F_{kl}-R^i_{\phantom{i}j}F_{kl}\right) \ ,
\end{equation}
where in the spherically symmetric case, the $b_1$ and $b_2$ terms vanish. Note that both variations yield gauge invariant expressions. Other terms may be added to $W_{EM}$, which would yield gauge dependant expressions. Nevertheless, some of these may still lead to gauge invariant equations of motion.

In the static spherically symmetric case, we substitute the ansatz \eqref{gSSS}, \eqref{ASSS} into the Lagrangian density to get
\begin{eqnarray}
\frac{\sqrt{f}}{N}r^6\mathcal{L}_{reduced}&\propto&3\LW r^8+2f'r^7+2fr^6-2r^6+\frac{1}{\LW}(2f'fr^5-2f'r^5-f^2r^4+2fr^4-r^4)+\nonumber\\*
&&{}-\frac{2f}{N^2}A'^2r^8+2p^2r^4+\frac{\beta p^2}{\LW}(2\LW r^4-2f'r^3+6fr^2-6r^2-5\beta p^2)+\nonumber\\*
&&{}+\frac{2f}{c^2N^2}(2N_r'N_rfr^7+N_r^2f'r^7+N_r^2fr^6) \ ,
\end{eqnarray}
where $\beta\equiv\mu^{-1}b_1$, we have set $\lambda=1$, and have omitted an overall factor independent or $r$.

We now find non-perturbative solutions with higher order magnetic terms. For the projectable case with $N_r\neq0$, $N=1$, and constant $f$, we find the solution:
\begin{equation}
\label{NrDB}
N_r=\pm\frac{c}{f}\sqrt{\frac{\LW}{2}r^2+f-1+\frac{2m}{r}+\frac{(f-1)^2}{2\LW r^2}-\frac{q^2+p^2}{r^2}+\beta p^2\left(-\frac{1}{r^2}+\frac{1-f}{\LW r^4}+\frac{\beta p^2}{2\LW r^6}\right)} \ .
\end{equation}
For the non-projectable case with $N_r=0$, we find
\begin{equation}
\label{N2DB}
N^2=f=-\LW r^2+1+\beta\frac{p^2}{r^2}-\sqrt{-4\LW mr+2\LW(q^2+p^2)} \ .
\end{equation}
These generalize the solutions with $\beta=0$ of \cite{charged,chargedprojectable}.

\section*{Acknowledgements}

I would like to thank Yaron Oz for his support and guidance.

\end{document}